# Agent A/B: Automated and Scalable Web A/B Testing with Interactive LLM Agents

### v3

@@ 5.2. Finding: Alignment with Human Behavior
Figure 6. Comparison of average customer spending across conditions, broken down by gender (a) and by age groups (b).:
5.3. Findings: System Effectiveness Across Interface Variants
To assess whether Agent A/B can detect subtle differences between design variants, we compared LLM agent behavior across control and treatment conditions featuring different filter configurations. As shown in Table 2, agents exhibited notable condition-dependent behavioral changes in response to the design manipulation.
Agents in the treatment condition clicked on more products (2.09 vs. 1.87), and the sessions have a higher total number of actions (6.60 vs. 6.05, t(998)=1.08, p=0.28). Most notably, the LLM agents in the treatment condition click filter options more frequently than the ones in the control condition (0.60 vs. 0.58, t(998)=1.00, p=0.32), which suggests that the treatment design that removes unrelated filters improved filter discoverability.
When looking at the final outcomes of the sessions in the treatment and control conditions, LLM-agent simulated virtual customers purchased more products in the treatment condition compared to the control condition(414 v.s. 404, $\chi^{2}(1)=5.51$, p=0.03). This represents a modest but statistically significant increase in purchase completion, suggesting that reducing the number of filter options improved the likelihood of customers making a purchase. Expressed in practical terms, the treatment condition resulted in roughly a 2–3% higher conversion rate, which could translate into millions of additional purchases per day on Amazon.
In terms of average spending across all customers, LLM agent simulated virtual customers in the treatment condition spent somewhat more than those in the control condition (\$60.99 v.s. \$55.14). However, this difference was not statistically significant (t(998) = 0.39, p = 0.69). The mean difference of about \$6 is small relative to the variability in spending, suggesting that while the treatment may have had a slight positive effect on purchase amounts, substantially larger sample sizes would be required to reliably detect such an effect.
Breaking down average spending by gender, the data show different patterns across groups. Female customers spent an average of \$49.18 in the control condition and \$51.30 in the treatment condition, indicating only a small increase under the treatment. Male customers, by contrast, showed a larger increase, from \$61.82 (control) to \$70.83 (treatment). While statistical tests were not conducted for each subgroup separately, the direction of change suggests that the simplified filter interface may have a stronger positive effect on male customers’ spending than on female customers.
A clearer set of differences emerges when comparing average spending across age groups. Among younger customers ($\leq 35$ ), average spending actually decreased in the treatment condition compared to the control condition ($\mathdollar 56.68\rightarrow\mathdollar 43.41$). Middle-aged customers (35–55) spent more under treatment ($\mathdollar 59.21\rightarrow\mathdollar 71.19$), while older customers (¿55) showed the largest increase ($\mathdollar 40.24\rightarrow\mathdollar 79.07$). These results suggest that the reduced filter interface may be particularly beneficial for older customers, potentially because it helps reduce choice overload. Conversely, younger customers may rely more on detailed filtering to find products that match their preferences, leading to lower spending when filter options are reduced.
In sum, reducing the number of filter options was associated with a higher purchase completion rate and possible increases in customer spending, though the latter effects did not reach statistical significance. Subgroup analyses suggest heterogeneity in response: males and older customers appear to benefit more from the simplified interface, while younger customers may prefer more filtering flexibility. Given the variability in spending outcomes, further testing with larger samples will be important to confirm these trends.
Notably, the LLM-agent A/B testing results aligned directionally with patterns observed in human A/B testing . This convergence increases confidence that agent-based simulations can anticipate real user behaviors and provide early insight into interface performance.
Overall, the A/B simulation conducted with Agent A/B shows that LLM-based agents, when guided by personas and embedded intentions, generate behaviorally rich and interface-sensitive interactions. These behaviors vary in predictable ways across personas, align with humans, and detect subtle differences in design. Such findings support the system’s role as a behaviorally grounded, scalable complement to conventional A/B testing. The results suggest Agent A/B is well-suited for early-stage prototyping, pre-deployment validation, and hypothesis-driven UX research.
6. Discussion
Our design and evaluation of Agent A/B demonstrate the feasibility and utility of using LLM-based autonomous agents at scale for behavioral simulation in the life-cycle of the web interface designs. These findings echo a growing trend in Machine Learning and HCI, where large-scale LLM agent-based simulations are increasingly explored as tools for social simulations, economic experiments, and iterative design support (Park et al., 2023; Horton, 2023; Gao et al., 2024). In this section, we discuss the broader implications of our findings, system-level affordance, current limitations, and future research directions.
6.1. Accelerating Design Iteration with Simulated Feedback
One of the most immediate and impactful contributions of Agent A/B lies in its ability to expedite the feedback loop for A/B testing and design evaluation in web-based applications. Traditional A/B testing, despite remaining the gold standard for user-centered and evidence-based design validation approaches, suffers from well-documented challenges that are primarily related to the experiment participants. In particular, real user traffic is often limited and expensive, feedback is slow due to the long time span for experiment execution, and there are not enough resources to test all the design variants (Fabijan et al., 2017; Kohavi et al., 2013; Xu et al., 2015). These structural constraints make it difficult for teams to validate novel ideas, particularly during early prototyping stages, when feedback is most critical. Furthermore, many practical challenges in engineering time and testing infrastructure often restrict the number of A/B testings that can be executed.
Recent work has emphasized the importance of early-stage, actionable feedback in early design phases to support interface and feature design (Fabijan et al., 2018). Tools like Apparition (Lasecki et al., 2015) and d.tools (Hartmann et al., 2006) have emphasized the value of rapid prototyping. However, most of these tools focus on interface prototyping rather than enabling feedback that is grounded in realistic user behavior.
Our Agent A/B, on the contrary, fills this critical gap by introducing a new phase into the design life-cycle: agent-based pilot experiments before real user testing. Our system enables a rapid and automated feedback loop through two key technical capabilities: 1) automatically scaling up massive amounts of LLM agents with realistic and diverse personas, and 2) asking the agents to perform realistic behaviors by “role-playing” real humans. These simulations are executed within live browser environments, with LLM agents continuously perceiving structured page representations, reason through goals and interface states, and predicting plausible next actions. The predicted actions are automatically translated into executions in web environments to picture individual user journeys. As a result, Agent A/B users, such as UX designers, can obtain large-scale, fine-grained insights into user decision strategies with rationales, failure cases, and behavioral divergence across design variants. These insights can support end users in iterating the designs and revising the actual user A/B tests before any real human participants are recruited or any feature is deployed.
6.2. Inclusive Piloting and Risk-Free Testing for Underrepresented Populations
In addition to accelerating iteration, Agent A/B enables another critical capability: inclusive, risk-free piloting for user populations that are otherwise difficult to recruit or ethically sensitive to test. These populations (e.g., older adults) are often underrepresented in early testing phases due to ethical or logistical barriers, and a poorly designed application may even impose harm on the user participants.
Agent A/B allows designers to simulate interactions from such groups by configuring agent personas that reflect the demographic and behavioral characteristics of targeted populations. For example, agents can be instantiated to simulate limited digital literacy, slower decision-making, or specific accessibility needs to reflect the behaviors of older adults more realistically and believably. By using our Agent A/B to generate a targeted agent population at scale and, more importantly, to automatically and iteratively test different design variants, designers can not only better evaluate how these groups might experience proposed features in real user testing but also fix design flaws and avoid any potential risks.
This capability not only enhances safety and inclusion in the design process but also ensures that interface and feature designs are more robust and accessible. Our idea echoes the early exploration of scaling interface evaluation using crowd-based approximations of end-user behavior (Bernstein et al., 2011; Oppenlaender et al., 2020). Our findings suggest that even in early prototyping phases, agents can offer intent-driven behavior patterns that may disclose usability bottlenecks, task friction, or potential failures. This paradigm shift reframes LLM agent simulation not merely as a performance benchmark but as a core design methodology. Agent-based piloting complements traditional user testing with risk-free experimentation, broadened design coverage, and prompt usability feedback that aligns with realistic interaction patterns.
6.3. Behavioral Fidelity of LLM Agents
Our results show that LLM-based agents can generate realistic behavior trajectories that are aligned with various human shopping strategies. The alignment in task completion rates, refinement behaviors, and interaction flow with real human data reflects the growing consensus in HCI that advanced LLMs can perform believable and context-aware behaviors (Park et al., 2023; Horton, 2023; Lu et al., 2025a).
However, simulated behavior is not a complete substitute for human cognition. As prior work in cognitive modeling and HCI simulation has emphasized (Gray and Altmann, 2001; John and Kieras, 1996; Ziebart et al., 2009), real humans are influenced by complicated explicit and implicit factors, including emotion, fatigue, prior experience, and latent goals. Many of these factors are not accessible to the current LLM agents. For instance, in our evaluations, agent behavior tended to be more deterministic, focused, and efficient compared to human users, who often displayed broader exploration and casual browsing. The behavior trajectories, as well as reasoning generated by LLM agents, should not be interpreted as perfect replications of user behaviors but as structured approximations that emphasize user behavior coverage and consistency. Our idea is analogous to the use of cognitive and computational models in prior HCI work, such as GOMS and ACT-R, which were not designed to replicate human behaviors perfectly but to estimate the interaction trade-offs at scale (Card, 2018; Olson and Olson, 1995; John and Kieras, 1996).
6.4. Toward Simulation-Supported Design Methodologies and Automated Design Optimization
While Agent A/B currently focuses on simulating A/B testing between interface design variants, our system architecture supports broader capabilities that move beyond this particular evaluation methodology. Our system laid the foundation for AI-based human behavior simulation at scale to support various design explorations, refinements, and evaluations. For instance, we can adapt our system to cognitive walkthrough and heuristic evaluation by designing the LLM agents to role-play different types of stakeholders (e.g., domain experts, HCI designers, technical developers, etc.) and specifying the agent’s intention to be aligned with the design methodologies correspondingly. This novel opportunity echoes a number of established works that underscore the development and adoption of automated tools for interface exploration and optimization promptly, such as Fuse (Kuznetsov et al., 2022), Apparition (Lasecki et al., 2015), and d.tools (Hartmann et al., 2006).
Further, our Agent A/B sheds light on the vision of “design mining” (Kumar et al., 2013), where large numbers of alternatives are evaluated computationally to identify the most promising solutions. Agent A/B supports this approach by serving as a simulation engine for a large pool of candidate variants and automating the design optimization through full interaction traces. For example, if $50$ layout variants of a filter panel were proposed, agent simulations could conduct large-scale simulations and rank the designs by search efficiency, user engagement, or failure rate across different user subgroups.
This extension of the system functions beyond as a testbed but as an interactive co-design partner to provide empirical evidence throughout the life-cycle of novel designs and functionalities. We view this direction as a natural evolution of the current system architecture and an exciting opportunity for future integration with generative design tools and adaptive interface frameworks.
6.5. Limitations and Future Directions
While Agent A/B establishes a strong foundation for the scalable and automated LLM agent-based A/B testing, we acknowledge two key limitations in the current scope of the system and inform ongoing development.
First, the robustness of our system and the behavioral fidelity of simulated agents remain bound by the reasoning and grounding capabilities of the underlying LLMs. Although recent LLMs such as Claude 3 and GPT-4o exhibit exceptional language understanding and planning capacity, they can still misinterpret complex or unconventional DOM structures, particularly in dynamically rendered web environments. Even state-of-the-art LLMs may fail to correctly parse web environments when encountering inconsistent element labeling, interactive latency, or unexpected modal interruptions (Zhou et al., 2023; Koh et al., 2024). While our system mitigates these issues through a novel implementation of leveraging JSON for structured environment parsing, LLM agent performance can still degrade when real-time content, along with the web environment structures, continuously updates and changes.
Second, agent behaviors in the current Agent A/B system are not able to comprehensively incorporate affective or meta-cognitive signals such as uncertainty, fatigue, or emotional response. These dimensions could be challenging for LLMs to model but are critical in shaping human cognitive behaviors and decision-making (Horton, 2023; Ziebart et al., 2009). LLM agents in our system tend to simulate goal-driven behavior with plausible cognitive structure but do not yet capture the full range of human variability or intent ambiguity.
Beyond these limitations, several technical and conceptual opportunities emerge for future work. One promising direction lies in expanding agent capabilities through multimodal information perception. For instance, integrating visual inputs (e.g., screenshots, spatial layout) with text-based content could allow agents to operate more robustly across richly designed or accessibility-diverse interfaces but still may introduce irrelevant information along the path. Recent progress in vision-language models and multimodal agents (Gupta and Kembhavi, 2023; Chezelles et al., 2024) suggests a pathway to generalizing agent behaviors beyond text-only interfaces.
Another important opportunity is collaborative or multi-agent simulation, especially for applications such as shared productivity tools, learning platforms, or social platforms. In these scenarios, agent interactions may involve negotiation, coordination, or conflict. Building on work like generative agents for social simulations (Park et al., 2023), future systems could simulate rich user ecosystems rather than independent task-solving to better reflect the collaborative work nature in real-world scenarios.
These directions demonstrate the extensibility of Agent A/B and its potential to evolve from a simulation platform into a broader foundation for behaviorally informed, user-model-driven design tools. As LLM-based interaction models continue to improve, we anticipate these agents will not only validate designs but also help co-create them.
7. Conclusion
In this paper, we presented Agent A/B, a system that enables large-scale, LLM agent-based simulation of A/B testing for web interfaces in real browser-based web environments. Our evaluation demonstrated that LLM agents exhibit realistic, goal-aligned behaviors, are sensitive to interface variations, and provide actionable feedback comparable to human users. By supporting rapid, risk-free behavioral piloting, Agent A/B introduces a new phase for agent-based piloting in the design life-cycle that complements traditional A/B testing and expands the scope of early-stage UX evaluation. We envision future extensions that further enhance agent fidelity, broaden domain coverage, and integrate simulation into intelligent design optimization workflows.
Appendix A Example of Persona Generated by Agent A/B
@@
[…]

### v4

@@ Appendix B A/B Testing Scenario
Figure 6. Comparison of average customer spending across conditions, broken down by gender (a) and by age groups (b).
[…]
Participant Gender Job Title Team
P1 F Project Manager PM Team
P2 M Software Development Manager Engineer Team
P3 F Product Manager PM Team
P4 M Product Manager PM Team
P5 M Product Manager PM Team
P6 M Machine Learning Researcher Science Team
Table 2. Demographics of formative study participants.
Appendix C Example of Persona Generated by Agent A/B
@@
Demographics:
Age: 35
Gender: Male
Education: Bachelors degree in Visual Communication
Profession: Freelance Graphic Designer
Income: \$70,000 (variable based on projects)
Financial Situation:
@@
Appendix D Persona Generation Prompt
You are a helpful assistant that generates diverse personas.
Examples:
[PERSONA EXAMPLE]
Generate a persona using the above examples. The persona should be different from previous personas to ensure diversity. The persona should:
- Have the age of {age}
- Be {gender}
- Have an income between \${income_range[0]} and \${income_range[1]}
Provide the persona in the same format as the examples.
Only output the persona, no other text.
Figure 7. The two design variants of the left filter panel on Amazon.com for the A/B testing case study with Agent A/B. In the control condition (left screenshot), all the filter options are shown; in the treatment condition (right screenshot), we apply a similarity-based ranking algorithm to reduce the filter options that have lower than 80% similarity to the user’s search query (e.g., “solar filter for telescope”).
A side-by-side comparison of two Amazon search results pages for the query ”solar filter for telescope.” Both pages show similar products but have differences in their filter options on the left sidebar. The left image displays additional filter categories under ”Department” such as ”Telescope Refractors,” ”Safety Goggles & Glasses,” and ”Telescope Eyepieces,” which are not present in the right image. Both images include nested subcategories under ”Camera & Photo Filters” and identical product listings on the right side of the page. This comparison highlights subtle differences in the user interface that could affect user navigation or filtering behavior.
Appendix E Interview Protocol
1. Overall Experimentation Practice
Can you describe your general workflow for running A/B testing experiments in your company?
2. End-to-End Workflow and Timing
How does an experiment progress from an initial design idea to a final deployment decision, and how long do the major stages typically take?
3. Pain Points and Challenges
What are the main challenges or friction points you encounter when designing, running experiments or analyzing results?
4. Experimentation Tools and Platform Evaluation
How well does the existing A/B testing experiment tool support your work, and what new features would you want?
5. Other
Is there anything else you would like to add, or do you have any questions for me?